OpenStudy (mrs.ambrose614):

Mrs.ambrose614 Human Calculator T Question 6 (Multiple Choice Worth 1 points) (05.05) Which of the following inequalities matches the graph? graph of an inequality with a dashed line through the points (0, 3) and (1, 9) with shading above the line −6x + y < 3 6x + y < 3 6x − y < −3 The correct inequality is not listed @satellite73 an hour agoEdit Question Share Question: @jim_thompson5910

myininaya (myininaya):

do you want to attempt to try to do this problem with me? if so have you found the slope yet?

OpenStudy (mrs.ambrose614):

No I have not

myininaya (myininaya):

you must calculate \[m=\frac{y_2-y_1}{x_2-x_1}\]

myininaya (myininaya):

where \[(x_1,y_1) \text{ and } (x_2,y_2) \text{ are given }\]

myininaya (myininaya):

try to plug in and tell me what you get for the slope

myininaya (myininaya):

\[m=\frac{9-3}{1-0}=?\]

OpenStudy (mrs.ambrose614):

4/1????

myininaya (myininaya):

9 take away 3 is 6

myininaya (myininaya):

6/1= ?

myininaya (myininaya):

any number divided by 1 is that number

OpenStudy (mrs.ambrose614):

6

myininaya (myininaya):

y=mx+b so the slope m is 6 y=6x+b you are given the y-intercept is (0,3) so you are given b

myininaya (myininaya):

so what is the equation of the line?

OpenStudy (mrs.ambrose614):

So the anwser is b???

myininaya (myininaya):

we haven't got to the answer yet

myininaya (myininaya):

you need to replace the b in y=6x+b

myininaya (myininaya):

the hint i gave is a bove

OpenStudy (mrs.ambrose614):

So b is 6??

myininaya (myininaya):

y=mx+b so the slope m is 6 y=6x+b you are given the y-intercept is (0,3) so you are given b b represents the y-intercept

myininaya (myininaya):

the y-intercept is at (0,3) so b is 3

myininaya (myininaya):

now shading above means you are going to have y greater

myininaya (myininaya):

so what do you think the inequality will be

OpenStudy (mrs.ambrose614):

So y is 6???

myininaya (myininaya):

no we just found m to be 6 and b to be 3 in the equation y=mx+b

OpenStudy (mrs.ambrose614):

So it's 3x????

myininaya (myininaya):

replace m with 6 and b with 3 in the equation y=mx+b do you know how to do that? try again

myininaya (myininaya):